At 12:14am UK local time, our monitoring reported an outage on two Vodafone provided fibre circuits delivered into our BOL8 location. These are both delivered over the same fibre bundle. These circuits provide our main office telephone service, and one of our metro ethernet fibre ring legs. After internal investigation… Continue reading →
A core switch failure in our U23 DC will have affected all connectivity briefly this morning, we are currrently investigating the cause, and are remedially migrating a couple of significantly affected hosts as a priority. A fuller RFO will follow.
We have just received an emergency maintenance notification from VirginMedia regarding one of our fibre optic metro ethernet links. Virgin Media will be performing emergency maintenance between midnight tonight (25/Sep/2022) and 8am. Our core network will automatically reroute during this window so no outages should be seen, however a reduction… Continue reading →
At 12:12pm today, our monitoring reported a failure of our VirginMedia metro ethernet leg within our metro ring. We have raised this as a fault with VirginMedia for investigation. We have a resilient ring so no outage or degration of service is in effect, only a reduction in resilience.
At approx 9:37pm tonight one of our SAN devices in BOL23 (EQL-SAS-4) reported that its primary controller was not responding. The controller triggered a kernel panic which caused the controller to reboot. The failover controller took over once it had confirmed the primary controller was not responding. This may have… Continue reading →
At 10:03am this morning, we saw a collection of BGP connections drop in Equinix MA1 (Manchester). These were spread across LINX MAN1, EquinixIX, and our GTT IP Transit connection. This caused a couple of minutes of instability whilst BGP reconverged and rerouted traffic automatically. A few minutes later, we saw… Continue reading →
We have identified a failure of one of our customer facing network switches in BOL8. Any customers connected directly to DS-173 in BOL8 will be affected. Other switches and customers are unaffected. We are investigating the cause and hope to have this switch back online shortly.
We had a brief network outage at approb 2:35pm today which was caused by a misconfiguration of our Vodafone fibre circuit due to human error. This affected any traffic traversing this link (such as our LINX LON1/LON2 traffic) and some traffic between Bolton and Manchester. This should now be resolved,… Continue reading →
This morning at approx. 6am UK local time, we had a switch failure of switch DS-194. This caused an outage for any directly connected customer. It also triggered an event in another switch which could affect connectivity to some virtual machines. At approx 8:45am, this switch was recovered and all… Continue reading →
We’re currently seeing one of our core fibre ring circuits offline (provided by Vodafone) between Manchester and Bolton. We’re not aware of any routine maintenance by Vodafone on this circuit so will proceed to investigate it as a fault. Due to our ring nature of the core, all traffic automatically… Continue reading →